find three numbers in which the second is 20 more than the first and the third is 5 less twice the first. their sum is 255.
Found 2 solutions by ankor@dixie-net.com, josgarithmetic:
Answer by ankor@dixie-net.com(22740)

find three numbers
a, b, c
in which the second is 20 more than the first
b = (a+20)
and the third is 5 less twice the first.
c = (2a-5)
their sum is 255.
a + b + c = 255
Replace b with (a+20); relace c with (2a-5)
a + (a+20) + (2a-5) = 255
combine like terms
a + a + 2a + 20 - 5 = 255
4a = 255 - 15
4a = 240
a = 240/4
a = 60
then
b = 60 + 20
b = 80
and
c = 2(60) - 5
c = 120 - 5
c = 115
:
You can ensure the sum of these 3 numbers are 255

Answer by josgarithmetic(39623)

FIRST              n
SECOND             n+20
THIRD              n-5
SUM                255